40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 18 22:35:23 2019, gautam, Update, SUS, ETMY actuator diagnosis  
    Reply  Fri Apr 19 11:36:23 2019, gautam, Update, SUS, No consistent solution for output matrix 
       Reply  Fri Apr 19 15:13:38 2019, rana, Update, SUS, No consistent solution for output matrix 
          Reply  Fri Apr 19 16:19:42 2019, gautam, Update, SUS, Actuation matrix still not orthogonal 
             Reply  Fri Apr 19 19:22:15 2019, rana, Update, SUS, Actuation matrix still not orthogonal 
    Reply  Mon Apr 22 22:43:15 2019, gautam, Update, SUS, ETMY sensor diagnosis  ETMY_sensorSpectra_consolidated.pdf
       Reply  Thu Apr 25 00:30:45 2019, gautam, Update, SUS, ETMY BR mode  ETMY_sensorSpectra_BRmode.pdfETMY_sensorSpectra_BRmode.pdfETMX_sensorSpectra_BRmode.pdfIMG_5993.JPG
Message ID: 14551     Entry time: Thu Apr 18 22:35:23 2019     Reply to this: 14554   14562
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY actuator diagnosis  

[rana, gautam]

Rana did a checkout of my story about oddness of the ETMY suspension. Today, we focused on the actuators - the goal was to find the correct coefficients on the 4 face coils that would result in diagonal actuation (i.e. if we actuate on PIT, it only truly moves the PIT DoF, as witnessed by the Oplev, and so on for the other DoFs). Here are the details:

  1. Ramp times for filter modules:
    • All the filter modules in the output matrix did not have ramp times set.
    • We used python, cdsutils and ezca to script the writing of a 3 second ramp to all the elements of the 5x6 output matrix.
    • The script lives at /opt/rtcds/caltech/c1/scripts/cds/addRampTimes.py, can be used to implement similar scripts to initialize large numbers of channels (limiters, ramp times etc).
  2. Bounce mode checkout:
    • ​The motivation here was to check if there is anomalously large coupling of the bounce mode to any of the other DoFs for ETMY relative to the other optics
    • The ITMs have a different (~15.9 Hz) bounce mode frequency compared to the ETMs (~16.2 Hz).
    • I hypothesize that this is because the ETMs were re-suspended in 2016 using new suspension wire.
    • We should check out specs of the wires, look for either thickness differences or alloying composition variation (Steve has already documented some of this in the elog linked above). Possibly also check out the bounce mode for a 250g load on the table top.
  3. Step responses for PIT and YAW
    • With the Oplevs disabled (but other local damping loops engaged), we applied a step of 100 DAC counts to the PIT and YAW DoFs from the realtime system (one at a time)
    • We saw significant cross-coupling of the YAW step coupling to PIT, at the level of 50%.
  4. OSEM coil coefficient balancing
    • I had done this a couple of months ago looking at the DC gain of the 1/f^2 pendulum response.
    • Rana suggested an alternate methodology 
      • we used the lock-in amplifier infrastructure on the SUS screens to drive a sine wave
      • Frequencies were chosen to be ~10.5 Hz and ~13.5 Hz, to be outside the Oplev loop bandwidth
      • Tests were done with the Oplev loop engaged. The Oplev error signal was used as a diagnostic to investigate the PIT/YAW cross coupling.
      • In the initial tests, we saw coupling at the 20% level. If the Oplev head is rotated by 0.05 rad relative to the "true" horizontal-vertical coordinate system, we'd expect 5% cross coupling. So this was already a red flag (i.e. it is hard to believe that Oplev QPD shenanigans are responsible for our observations). We decided to re-diagonalize the actuation.
      • The output matrix elements for the lock-in-amplifier oscillator signals were adjusted by adding some amount of YAW to the PIT elements (script lives at /opt/rtcds/caltech/c1/scripts/SUS/stepOutMat.py), and vice versa, and we tried to reduce the height of the cross-coupled peaks (viewed on DTT using exponential weighting, 4 avgs, 0.1 Hz BW - note that the DTT cursor menu has a peak find option!). DTT Template saved at /users/Templates/SUS/ETMY-actDiag.xml
      • This worked really well for minimizing PIT response while driving YAW, not as well for minimizing YAW in PIT. 
      • Next, we added some YAW to a POS drive to minimize the any signal at this drive frequency in the Oplev YAW error signal. Once that was done, we minimized the peak in the Oplev PIT error signal by adding some amount of PIT actuation.
      • So we now have matrices that minimize the cross coupling between these DoFs - the idea is to back out the actuation coefficients for the 4 OSEM coils that gives us the most diagonal actuation, at least at AC. 
  5. Next steps:
    • All of our tests tonight were at AC - once the coil balancing has been done at AC, we have to check the cross coupling at DC. If everything is working correctly, the response should also be fairly well decoupled at DC, but if not, we have to come up with a hypothesis as to why the AC and DC responses are different.
    • Can we gain any additional info from driving the pringle mode and minimizing it in the Oplev error signals? Or is the problem overconstrained?
    • After the output matrix diagonalization is done, drive the optic in POS, PIT and YAW, and construct the input matrix this way (i.e. transfer function), as an alternative to the usual free-swinging ringdown method. Look at what kind of an input matrix we get.
    • Repeat the free-swinging ringdown with the ETMY bias voltage adjusted such that all the OSEM PDmons report ~100 um different position from the "nominal" position (i.e. when the Y arm cavity is aligned). Investigate whether the resulting eigenmode frequencies / Qs are radically different. I'm setting the optic free-swinging on my way out tonight. Optic kicked at 1239690286.
ELOG V3.1.3-